OpenStudy (anonymous):

Two buses leave Houston at the same time and travel in opposite directions one bus averages 55mph, the other bus averages 45mph when will they be 400 miles apart

OpenStudy (zzr0ck3r):

\(55x+45x=400\)

OpenStudy (anonymous):

relative speed=55+45=90 mph distance=400 miles time=distance/speed